OpenStudy (anonymous):

What are the coordinates of the vertex of the graph of the function y = –x² + 6x –11? (3, –2) (3, 16) (–3, –29) (–3, –20)

OpenStudy (zehanz):

Hey @iloveyooh: this is your lucky day! The x-coordinate of the vertex is the same as the x-value of the line of symmetry (see your previous question), So x=-b/(2a). Calculate y by putting the x you found in the formula y=-x²+6x-11.

OpenStudy (anonymous):

could you write the problem out? because I don't understand how to put what you just said into a problem.

OpenStudy (anonymous):

@ZeHanz

OpenStudy (anonymous):

I NEED HELP!!!!!!!!!!

OpenStudy (zehanz):

Well, if x=-b/2a, and a=-1, b=6, I don't see what could be confusing about calculating this...